OpenStudy (anonymous):

help understanding this answer:

OpenStudy (ja1):

sure, go ahead.

OpenStudy (anonymous):

\[\int\limits_{0}^{4} \left| \sqrt(x)-1 \right|=2\] Why?

OpenStudy (ja1):

um, to advanced for me lol sorry ask someone more advanced.

OpenStudy (callisto):

For \(0\le x < 1\) , \(\sqrt{x} -1\) <0 For \(1\le x \le 1\) , \(\sqrt{x} -1 \ge\)0 So, \[\int\limits_{0}^{4} \left| \sqrt(x)-1 \right|dx\]\[=\int_0^1 (1-\sqrt{x}) dx + \int_1^4 (\sqrt{x}-1)dx\]Can you do it from here?

OpenStudy (anonymous):

@callisto \[\int\limits_{0}^{1} [x-\frac{ 2 }{ 3 }x^\frac{ 3 }{ 2 }] +\int\limits_{1}^{4} [\frac{ 2 }{ 3 }x^\frac{ 3 }{ 2 }-x]= [1/3+16/3-5/3]=4\] But the answer supposedly is 2. Is this wrong?

OpenStudy (callisto):

It is wrong. You made some mistakes?!

OpenStudy (anonymous):

Can you help me with it till the end? I can't seem to figure it out...

OpenStudy (callisto):

\[\int\limits_{0}^{1} [x-\frac{ 2 }{ 3 }x^\frac{ 3 }{ 2 }] +\int\limits_{1}^{4} [\frac{ 2 }{ 3 }x^\frac{ 3 }{ 2 }-x]\]\[= [1-\frac{2}{3}+\frac{16}{3}-4 - \frac{2}{3}+1]\]\[= [1-\frac{4}{3}+\frac{16}{3}-4+1]\]\[= [1+\frac{12}{3}-4+1]\]\[= [1+4-4+1]\]\[=2\]

OpenStudy (callisto):

The first line should be \[ [x-\frac{ 2 }{ 3 }x^\frac{ 3 }{ 2 }]_0^1 +[\frac{ 2 }{ 3 }x^\frac{ 3 }{ 2 }-x]_1^4\]

OpenStudy (anonymous):

thanks! You're amazing! Greatly appreciated it!

OpenStudy (callisto):

You're welcome :)